Suppose that a death body is found inside a room that is kept at a constant 65 F. The body's temperature was first measured to be 800 F at 7:00 am, and an hour later 785' F. Assume that the person was healthy before she/he died and had a body temperature of 98.6 F.Use the Newton's Law of Cooling Model to determine the time when the person died. (a) Write the corresponding initial value problem describing the situation (b) Find the proportionality constant (c) Find the integration constant
at your solution satisfies the DEQ and the initial condition (d) Check th (e) Find an estimate of the time (hour and minute) when the person died.
Answer #1

coolins i Acco dins to Neutons Jau df 7Ooarm) altt (b) Rearra ny kolt 7-65 -) T= 65+e 13-3 e. -)T= 65 +15e nital condtion a

The person died at 11:21 pm previous night.

Comment if any doubt.